Nash Inc. issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $121,000.

(a) Prepare the journal entry for the issuance when the market price of the common shares is $164 each and market price of the preferred is $205 each.
(b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $212 per share.

(Round answers to 0 decimal places, e.g. $1,225. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1

Answer

A. journal entry for the issuance when the market price of the common shares is $164 each and market price of the preferred is $205 each.

in the question it has given two situations

Thus we have to calculate the total amount of capital and common and preffered capital seperately. It can be calculated by,

Fair value of common share = 500 * 164 = 82000

Fair value of preffered share = 100 * 205 = 20500

Total 102500

Then the total capital have to be allocated to common stock and preffered capital.It can be done by,

Allocation to common stocks

82000/102500 * 121000 = 96800

Allocation to Preffered stocks

20500/102500 * 121000 = 24200

Total Allocation

96800+24200 = 121000

Now we have calculated the amount of capitals in common stock and preffered capital, thus we can enter the journal entries for thay as follows.

Journal Entries

Cash A/C Dr. 121000

To Common Stock A/C 5000

To Paid instock capital excess of par A/C    91800 (96800 - 5000)

To Preffered Stock A/C     10000

To Paid in pref. capital excess of par A/C   14200 (24200 - 1000)

B.journal entry for the issuance when only the market price of the common stock is known and it is $212 per share.

Lumpsum amount 121000

Allocated to common shares 106000 (500 * 212)

Balance allocated to preference shares 15000

Now we have calculated the amount of capitals in common stock and preffered capital, thus we can enter the journal entries for thay as follows.

Journal Entries

Cash A/C Dr. 121000

To Common Stock A/C 5000

To Paid in capital excess of par A/C    101000 (106000 - 5000)

To Preffered Stock A/C     10000

To Paid in capital excess of par   A/C 5000 (15000 - 10000)